OpenStudy (anonymous):

What is the radius of an expanding circle at a moment when the rate of change of its area is numerically twice as large as the rate of change of its radius?

OpenStudy (anonymous):

*

OpenStudy (anonymous):

A= pir^2 dA/dt = 2pir dr/dt (dA/dt)/(dr/dt) = 2 = 2pir r=1/pi